[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Aug 19, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Aug 19 20:09:56 UTC 2017


HSU Shift Log for Sat Aug 19 2017 16:09:50 GMT-0400 (EDT)

Shift: H

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: After GOES-16 meso sector for Harvey expires at 21Z  
today, there will be no more requests honored for 29 hours because of  
5-minute full disk testing.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170819/6db7b6fd/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list